OpenStudy (anonymous):

solve for x 2(x+8)+5(x-1)=6

OpenStudy (anonymous):

i solved the whole thing,just couldnt get the right answer..

OpenStudy (anonymous):

2(x+8) + 5(x-1) = 6 2x + 16 + 5x - 5 = 6 2x + 5x = 6 -16 + 5 7x = -5 x = -5/7

OpenStudy (anonymous):

thanks,appreciate it..

OpenStudy (anonymous):

You're welcome.